OpenStudy (anonymous):

solve the equation 2/3a=8

OpenStudy (anonymous):

2/3a=8 2=8(3a) 2/24=(a)=1/12

OpenStudy (anonymous):

(2/3)a=8 a=8(3/2)=12

OpenStudy (anonymous):

Is it: \[\frac{2}{3}a=8\] or \[\frac{2}{3a}=8\]

OpenStudy (anonymous):

yeah w/c one is it? any way its there w/c one supposed to be.. lol....

OpenStudy (anonymous):

its 2 over 3 a not 2 over 3a

OpenStudy (anonymous):

Im not sure how to type it the way rudy did. :/

OpenStudy (anonymous):

Well, mark o. covered both options, so just pick the one you need. Also, for typing equations, press the equation button right under box where you type.

OpenStudy (anonymous):

you need to have a math software installed to you pc,,,i cant installed them myself cause my C is full...lol

OpenStudy (anonymous):

Oh ok sweet! I see. Thanks

OpenStudy (anonymous):

The other way is tricky, but try this: \[\frac{2}{3}a\&] Type the line above but eliminate the ampersand (The &)

OpenStudy (anonymous):

\[\frac{2}{3}a\]

OpenStudy (anonymous):

wow great thnx rudy i learned something..lol

OpenStudy (anonymous):

I got that from the equation editor, then if you use them enough you'll memorize them and it can be easier.

OpenStudy (anonymous):

alright thnx rudy you ARE A BIG HELP

OpenStudy (anonymous):

god luck now guys,,have fun.thnx ....lol